OpenStudy (sally):

In 1967 car valued at $800; in 2006 valued at $27000. find the exponential growth? I think it is 1/39 In 27000/800

OpenStudy (apples):

Correct, \[27000 = 800e^{39r}\]\[\frac{27000}{800} = e^{39r}\]\[\log{\frac{27000}{800}} = 39r\]\[\frac{1}{39} \log{\frac{27000}{800}} = r\]

OpenStudy (sally):

with the same information as above i have 2 questions. a. Estimate teh value of the car in 2008? b. After how long will the value of the car be $40,000 (assuming that there is no change in the growth rate.